To gauge their fear of going to a dentist, a random sample of adults completed the Modified Dental Anxiety Scale (MADA) questionnaire. Scores on a scale range from zero (no anxiety) to 25 (extreme anxiety). The mean score was 11 and the standard deviation was 4, i.e. a normal distribution with 𝜇=11 and 𝜎=4. Find the probability that someone scores between 10 and 15 on the MADA (by hand and using MS Excel).

Find the probability that someone scores above 20 on the MADA


1
Expert's answer
2021-10-12T10:26:43-0400

"\mu=11 \\

\sigma= 4"

Find the probability that someone scores between 10 and 15 on the MADA

"P(10<X<15) = P(X<15) -P(X<10) \\

= P(Z< \frac{15-11}{4}) -P(Z< \frac{10-11}{4}) \\

=P(Z< 1) -P(Z< -0.25) \\

= 0.8413 -0.4013 \\

= 0.4400"

Find the probability that someone scores above 20 on the MADA

"P(X>20) = 1 -P(X<20) \\

= 1 -P(Z< \frac{20-11}{4}) \\

= 1 -P(Z< 2.25) \\

= 1 -0.9877 \\

= 0.0123"